OpenStudy (anonymous):

Find the series shown (will be typed below) A. 7 + 9 + 11 + 13 + 15 B. 11 + 13 + 15 + 17 + 19 C. 11 + 13 + 15 + 17 + ... D. 7 + 9 + 11 + 13 + ...

OpenStudy (anonymous):

\[\sum_{n=3}^{\infty}2n+5\]

ganeshie8 (ganeshie8):

put n = 3 first term = 2(3) + 5 = 11

ganeshie8 (ganeshie8):

so its either B or C

ganeshie8 (ganeshie8):

since its an infinite series, it never stops

ganeshie8 (ganeshie8):

so its C

OpenStudy (anonymous):

Thank you so much!

ganeshie8 (ganeshie8):

yw :))